OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4.

OpenStudy (anonymous):

y-y1=slope(x-x1) where( -3, -1) =(x1 y1)

OpenStudy (anonymous):

i dont understand how to do the formula

OpenStudy (anonymous):

1) you have to use point slope formula which is y-y1=m(x-x1) 2) your y1 is -1 and your x2 is -3 3) your equation will look like y+1=4(x+3) 4) then use the distributive property to get y+1=4x+12 5) subtract 1 from both sides to get y=4x+11 6) that is your answer Hope that helps :D

OpenStudy (anonymous):

oh that makes alot more sence. thank you so much. now i will understand how to do the rest :)